Mustang Enterprises, Inc., has been considering the purchase of a new manufacturing facility for $280,000. The facility is to be fully depreciated on a straight-line basis over seven years. It is expected to have no resale value after the seven years. Operating revenues from the facility are expected to be $115,000, in nominal terms, at the end of the first year. The revenues are expected to increase at the inflation rate of 2 percent. Production costs at the end of the first year will be $40,000, in nominal terms, and they are expected to increase at 3 percent per year. The real discount rate is 5 percent. The corporate tax rate is 40 percent.
   
Calculate the NPV of the project. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

FYI $78,900.50 IS NOT THE ANSWER!!!
Answer #1

calculations

straight line depreciation = (cost of the facility- salvage value)/life of the facility

=(280,000-0)/7= $40,000

year 1 2 3 4 5 6 7
revenue inflows 115,000 117,300 119,646 122,039 124,480 126,970 129,510
revenue outflows(-) 40,000 41,200 42,436 43,710 45,021 46,371 47,762
net cash flows 75,000 76,100 77,210 78,329 79,459 80,599 81,748
depreciation(-) 40,000 40,000 40,000 40,000 40,000 40,000 40,000
income before taxes 35,000 36,100 37,210 38,329 39,459 40,599 41,748
Taxes@40% (-) 14,000 14,440 14,884 15,331 15,783 16,239 16,699
after-tax income 21,000 21,660 22,326 22,998 23,676 24,360 25,049
depreciation (+) 40,000 40,000 40,000 40,000 40,000 40,000 40,000
after-tax cash flows 61,000 61,660 62,326 62,998 63,676 64,360 65,049
after-tax sal. value 0 0 0 0 0 0 0
after tax net cash flow 61,000 61,660 62,326 62,998 63,676 64,360 65,049
discount rate @ 5%(*) 0.952 0.907 0.864 0.823 0.784 0.746 0.711
PV of cash flows 58,072 55,925.62 53,849.66 51,847.35 49,924.33 48,012.56 46,249.83
total NPV 363,881.35

after tax salvage value= cash proceeds - tax on gain or loss

tax on gain or loss= (cash proceeds - book value)* tax rate

OR after tax salvage value= cash proceeds -(cash proceeds- book value)* tax rate

after tax salvage value is zero as according to question there is no resale value after 7 years

# note discount rate =(1+r)-n where r is rate of discounting and n is no. of period

= (1+0.05)^-1= 0.952

=(1+0.05)^-2=0.907

so on